# AppSupplemental Materials References

### v1

@@ Abstract
The United States leads the world in the number of violent mass shootings that occur each year, and policy making on firearms remains polarized along party lines. Are legislators responsive to mass shootings? We estimate the latent positions of nearly 2,000 state legislators on gun policy from their roll-call voting records on firearm-related bills from 2011 to 2022. Employing a staggered difference-in-differences design, we find that mass shootings within or near a state legislator’s district do not alter their voting behavior on firearm policy, on average, for members of both parties. Our estimated effects of mass shootings on treated legislators’ support for restrictive gun policies (on a -1 to 1 scale) range from a 4.8% reduction among California Democrats and a 0.9% increase among California Republicans to, across six total states, a 5% (among Democrats) and 7.1% (among Republicans) increase, with 95% confidence intervals spanning opposite directions. We conclude that, on average, mass shootings fail to produce changes in a legislator’s support (opposition) for restrictive (permissive) firearms bills. Our findings suggest that even the most heinous acts of mass violence — that are squarely in the domain of events that state legislators might respond to — fail to produce any measurable effects on legislators’ positions on firearm-related policy.
@@ 1 Introduction
The current scholarship on behavioral responses to mass shootings — an example of a crisis event that legislators might respond to — has particularly focused on the effects of these events on public opinion (Gunn et al., 2018; Barney and Schaffner, 2019; Newman and Hartman, 2019; Rogowski and Tucker, 2019), the electoral behavior of affected constituencies (Hassell, Holbein and Baldwin, 2020), and the firearm-related policy environments of states (Luca, Malhotra and Poliquin, 2020), but the results are decidedly mixed. In some instances, scholars find that mass shootings cause surges in engagement with gun policy (Reny et al., 2023) while others find that other responses are limited by temporal, spatial, and political factors (Sharkey and Shen, 2021). In a similar vein, it appears that some mass shootings — particularly those with high levels of media coverage — coincide with increases in firearms purchasing while others do not (Iwama and McDevitt, 2021; Liu and Wiebe, 2019). Additional evidence suggests that legislators may lack the ability or incentive to adjust their behavior even when facing possible ideological, environmental, or electoral pressures (Butler and Hassell, 2018; Bromley-Trujillo and Poe, 2020). Even if we believe that mass shootings might lead to shifts in opinions on firearm-related policies — which is not obvious, as per Baxter-King (2024) — there may not be enough of a change in the electoral behavior of affected constituencies to cause a legislator to alter their voting behavior on relevant policies, even as some scholars find that members of Congress who are facing reelection pressures, and represent sizable “pro-gun constituencies”, may be more likely change how they vote on gun-related legislation (Bouton et al., 2014). Furthermore, it is not clear that changes in the voting behaviors or policy preferences of a legislator’s constituency (should they occur) would prompt changes in a representative’s actions (Rogers, 2017).
@@
Our paper investigates the effects of exposure to mass shootings on the voting behavior of state legislators who represent the communities most directly impacted by these events. We focus our analysis on estimating the aggregate effect of mass shootings on the behavior of lawmakers rather than shooting-specific effects. We combine roll-call voting records of California state legislators with a novel dataset from the Giffords Law Center (GLC) that ranks every firearm-related bill introduced in the California State Assembly and Senate from 2011 to 2022 on a permissive to restrictive scale. [^4] We validated the GLC’s bill rankings, which is discussed in greater detail in S.3. Using pairwise comparison and the Bradley-Terry model (Bradley and Terry, 1952), we construct a yearly score of each individual lawmaker’s directionality of support for gun control legislation from the 2011-2012 to 2021-2022 sessions. Our scoring regime covers 264 individual California legislators over twelve years.
@@
Using this measure of an individual legislator’s latent position on firearms policy over time, we identify which legislators represented districts that experienced a mass shooting between 2011 and 2022. Through a difference-in-differences design accounting for heterogeneity in treatment timing, we estimate whether the occurrence of a mass shooting within a legislator’s district causes a measurable change in their position on gun policy. We find that when a mass shooting occurs within a legislator’s district, California lawmakers reduce their support for gun control policies by approximately 1.45% compared to all other legislators (ATT = -0.029, SE = 0.062, outcome measured on a -1 to 1 scale), though this estimate is a precise and statistically insignificant null. Our results find no evidence of such responsiveness and these statistically insignificant estimates persist across multiple estimation strategies and varied treatment assignments (such as estimations accounting for the distance of a legislator’s district from a mass shooting). This finding conflicts with the self-reported assessments of responsiveness to mass shootings that these same legislators described in a series of qualitative interviews we conducted. In sum, our mixed-methods approach in California finds that legislators do not change their voting behavior following a mass shooting — though this finding only speaks to a single metric of legislator behavior in a single state.
@@
As a robustness check on our main findings for California, we analyze the effects of mass shootings on the gun control positions of state legislators in five additional states: Colorado, Florida, Georgia, Missouri, and Texas. [^5] The GLC ranks every state in terms of how restrictive the gun policy environment is. In this paper, we focus on states that had at least one mass shooting occur during our period of study and select one state from each of the quintiles of the GLC ranking: California (1), Colorado (10), Florida (24), Georgia (33), and Missouri (48). Texas (31) is also included as it experienced a large number of shootings (6) during our period of study, and has a more permissive policy environment than California. This offers an opportunity to assess the effects of mass shootings in varying firearms policy environments. These states all experienced at least one mass shooting during the time period we study in this paper. Ultimately, we find that across a twelve year period, 28 mass shootings — occurring in 6 states and requiring the scoring of nearly 2,000 individual state legislators over time — caused no discernible change in the gun control positions of state legislators on regulatory proposals on firearms, on average, when such an event occurs within their districts. We also inspect other forms of behavior that legislators may engage in: similar to our main results, we find that legislators do not cosponsor or author more or fewer firearm-related bills following a mass shooting, on average.
@@ 2 Theory and Qualitative Findings
To better inform our expectations regarding the effects of mass shootings on the voting behaviors of state legislators, we conducted a series of qualitative interviews with five current and former members of the California state legislature who represented districts in which a mass shooting had occurred between July 2010 and 2022.666Additional details on the procedures and results of these qualitative interviews can be found in section S.1 of the Supplemental Materials. In this section, we introduce the literature that informs our theoretical expectations, and discuss the results of these qualitative interviews that support and contradict our priors.
@@ 2.1 Do Mass Shootings Incentivize Agenda Setting or Act as Windows of Policy Opportunity?
Political agenda setting — which can be conceptualized as political elites’ ability to influence public opinion and public policy, for instance — has long been considered the “prime instrument of power” wielded by lawmakers to alter the nature of political conflict (Schattschneider, 1960). At the sub-national level, political agenda setting is largely driven by questions of issue prioritization, as state or local officials determine which areas to focus their rather limited policy making resources on addressing (Franklin and Ebdon, 2004; Erie, Kogan and MacKenzie, 2011). Mass shootings, as highly salient events that capture the attention of the mass public and media, may then serve as an opportunity for state legislators to set legislative priorities and build on the public’s attention towards specific issues. As Legislator A (D, Southern California) said in their interview with us, the shooting in their district offered them an opportunity to “use the moment, in all its horrifying nature, to also give [the victim’s] lives and their deaths meaning.”
@@ 2.2 The Limited Effects of Mass Shootings in an Era of Partisan Polarization
The results from our qualitative interviews suggest that partisan polarization may play a substantial role in informing the responsiveness of legislators to mass shootings that occur within their respective districts. Legislator E, a Democrat, spoke about how partisan polarization certainly informed their responses to the shooting in their district, saying “[I was] trying to walk that very fine line because of political differences in a community that’s more conservative. There were definitely some families [of victims] that were a little more Second Amendment oriented than others.” Legislator C, a Republican, spoke about the role of partisan polarization more bluntly, explaining what their party told them after the shooting: “I was told directly, if I speak about [gun control], it’s going to cost \$1,000,000 per Tweet. For us to defend you in the next election, it’s going to cost 1,000,000 bucks a Tweet.” The direct threat of sanctions from party leadership that would affect a legislator’s campaign warchest suggests that any sort of defection — even one as mundane as a Tweet — from policy positions preferred by the party would result in severe consequences. [^8] We cannot confirm the veracity of Legislator C’s claims about sanctions threatened by the California GOP. We can, however, confirm that Legislator C did not post a single Tweet related to a firearms policy proposal until 181 days after the shooting. This Tweet was expressing opposition to a permissive policy.
@@ 3.1 Mass Shootings in Scope of Study
Note: Red districts are those affected by a mass shooting, shown as the small, light red points at the location of the shooting. The map uses district boundaries adopted after the redistricting cycle following the 2010 Census, and only plots shootings that affected districts under these new maps (accounting for variation in timing of adoption of these maps across states). For a full list of affected districts and shootings (including those happening prior to redistricting), see S.2. Nebraska has a unicameral state legislature, so we do not report lower chamber districts and only show upper chamber districts on the map.
Additional details on these shootings and the selection decisions underlying the inclusion of these states in our analyses can be found in Table S.2.1. For California, we consider the effects of eleven mass shootings.
@@ 3.2.1 GLC Firearm Bill Rankings and Bill Effects
From 2011 to 2023, the GLC hand-coded every bill introduced or enacted by every state legislature that was directly related to firearms policy — with the universe of legislation ranging from provisions allowing the permitless carry of concealed firearms, to large-scale budgetary items mentioning firearms in a single sub-item, to bans on assault-style weapons. This yields a total of 15,299 legislative attempts at the state level. For each of them, the GLC ranked it on a “weaken/neutral/strengthen” scale with regards to the bill’s likely effect on a state’s respective gun control policy environment (with “strengthen” bills making a state policy more restrictive and therefore more favorable to gun violence prevention advocates, for instance). [^10] Under the terms of a non-disclosure agreement dictating the amount and type of data that we are able to share in replication materials and publications, we arranged for conditional access to this as-of-yet untapped source of internal lobbying data. In Figure S.3.1, we plot the number of bills introduced in California over each legislative session, according to the GLC Firearm Bill Rankings, and find that California saw a large increase in the number of restrictive bills introduced over the period we study.
@@
In section S.3 of the supplemental materials, we discuss our validation procedure used for the GLC Bill Rankings, and we ultimately accept that “weaken” laws (according to the GLC) can generally be considered permissive (-1), “neutral” as neutral (0), and “strengthen” laws as restrictive (+1). We refer to this -1/0/+1 score as the “Bill Effect” throughout this paper — a novel measure of the relative restrictiveness or permissiveness of a firearm-related legislative measure introduced in a state. Table 1 shows three examples of bills from California during the period that we study, with the original GLC ranking and the Bill Effect.
@@ 3.2.2 Measuring Legislators’ Latent Position on Firearm-related Issues
Assuming that each pairwise comparison is independent, the latent trait $\lambda$ can be estimated by maximum likelihood. As we wish to compare legislators across time, we utilize a “bridge” legislator for every state who served in all sessions for which we have data — in California, Tori Atkins is used as our anchor to align the latent scores ($\lambda_{i}$). [^12] The assumption is that the bridging legislators’ underlying ideology changes only modestly from one year to the next. We select Toni Atkins because she is never treated in any of the specifications and she was a consistent leader in Democratic politics in CA during our period of study, serving as President Pro Tempore of the California State Senate, and previously as the Majority Leader of the California State Assembly. This assumption does not imply cross-state comparability. See Bailey (2007); Lewis and Tausanovitch (2015); Lewis and Sonnet (2020) for discussions on the limitations on the comparability of ideal point estimates. Finally, we rescale the scores to range from -1 to 1, from least to most supportive of restrictive firearms policies. For California legislators, the Gun Control Scores are calculated using 1927 votes on firearm-related issues (on average, 160.6 votes each year).
Unlike the conventional methods of using a spatial choice model or Bayesian Item Response Theory (IRT) model on a subset of roll-call voting records, our scoring regime uses the direct Bill Effects to orient scores in the direction of supporting gun control policies. As robustness checks, we also estimate latent traits using the IRT model (Clinton, Jackman and Rivers, 2004) in S.5.5 and discuss alternative measures such as survey responses in S.4.1. Figure 2 shows the distribution of the Gun Control Scores over time for California state legislators. We pick up a small, upward trend for all California state legislators over the years, indicating that California’s gun policy environment has become more restrictive just as the majority of legislators have become more favorable to restrictive policies, on average. However, there is still a clear separation between Democratic and Republican legislators, suggesting that partisan legislators are polarized on this policy domain.
@@ 3.3 Treatment Assignment and Causal Identification
As seen in Figure 1 and Table S.2.1, the shootings we consider as “treatments” in this paper happened in different legislative districts across a wide time period, and in different states. Given this heterogeneity in treatment timing and location, traditional difference-in-differences (DiD) estimators (such as two-way fixed effects) are not appropriate for this setting as they fail to account for fluctuations in the effects of treatment due to adaptation, changing environmental conditions, or the immediacy of responses among treated units (Goodman-Bacon, 2021). For these reasons, we follow the staggered DiD estimation strategy proposed by Callaway and Sant’Anna (2021). In Section S.5.2, we validate these results by employing a stacked DiD estimation strategy similar to one used by Cengiz et al. (2019) to account for varied treatment timing.
@@
The two central assumptions underlying these approaches are that of staggered treatment adoption and parallel trends with never-treated units. For this first assumption, we posit that once units (legislators in state A) receive a treatment (a shooting in state A occurring within or near their district, depending on the specification), they remain treated throughout the entire period of observation and treatment is irreversible. For parallel trends with never-treated units, we assume that groups of units during periods where treatment is not yet applied serve as control units — in other words, legislators who have not had a mass shooting occur within or near their district (and never do within the time frame we study) serve as controls — and these units would have followed similar trends in voting behavior in all post-treatment periods. This assumption relies on two conditions that our setting satisfies: (1) there must be a sufficiently powered group of units that never received treatment (we have 240 never-treated legislators in California, for example), and (2) these never treated units must be similar to those that receive treatment so that we can compare their outcomes. Additionally, we assume that there is limited anticipation of the treatment among all groups of units that are eventually treated. We report our validation and sensitivity analysis for these assumptions in Section S.5.1.
3.3.1 Formal Models
$ATT(g,t)=\mathbb{E}[Y_{t}(g)-Y_{t}(0)|G_{g}=1]$ (2)
where $g$ denotes a particular group of treated units at time period $t$. Hence, $ATT(g,t)$ is the expected difference between the observed Gun Control Score for treated legislators and untreated legislators at time t. In other words, this yields the group-time average treatment effect (Callaway and Sant’Anna, 2021; Baker, Larcker and Wang, 2022). In this setting, we consider a legislator as treated if they represented a district in which a shooting occurred at the time of the shooting. In this sense, we consider the effects of a shooting within legislators, and not within districts. This is detailed in the following section.
where $y_{it}$ denotes the Gun Control Score for legislator $i$ in time $t$ and group $g$. $\alpha_{2}^{g,t}$ represents the effect associated with group g at time t, with $\mathbb{I}[E_{i}=g]$ indicating if legislator i belongs to group g. We also include a time indicator effect, $\alpha_{3}^{g,t}\times\mathbb{I}[T=t]$, where $\alpha_{3}^{g,t}$ represents the effect associated with time t for group g. Using never-treated legislators as our control units, $\beta^{g,t}$ serves as an estimator for $ATT(g,t)$, as this term represents the interaction effect for group g at time t. In the formal specification of our model for California, we also include interacted legislator- and chamber-fixed effects. In models where we expand our study to include additional states, we include state-fixed effects with the interacted legislator- and chamber-fixed effects.
Ultimately, the unbalanced nature of our panel poses some issues for estimation insofar as we are limited in our ability to explicitly model the outcome evolution for individual legislators who drop out of (and, sometimes, back into) the data. Therefore, we use inverse probability weighting to model the conditional probability of legislator i being in group g, or an appropriate control group, given a set of included covariates. [^13] In the case of models for only California lawmakers, these covariates are unique legislator identifiers and an indicator for chamber, and for models including multiple states, we also include a state indicator. The benefit with this approach is that we essentially inflate the weights for units who are potentially under-represented due to the imbalanced nature of our data — an issue that is inherently present with a panel of legislators who rarely remain in office for the entire period of study.

### v2

# AppAppendix References

@@ Abstract
The United States leads the world in the number of mass shootings that occur each year even as policy making on firearms remains polarized along party lines. Are legislators responsive to mass shootings? We estimate the latent positions of 264 California state legislators on gun policy from their roll-call voting records on firearm-related bills from 2011 to 2022. Employing a stacked difference-in-differences design, we find that mass shootings within or near a state legislator’s district do not alter their voting behavior on firearm policy, on average (with 95% confidence intervals spanning opposite directions). When we extend our analyses to an additional 5 states (for a total of 1,864 legislators), we find similarly null results, though we report evidence of statistically significant effects in both directions from event-specific estimates. Our findings suggest that even the most heinous acts of mass violence generally fail to produce measurable effects on legislators’ positions on firearm-related policy.
@@ 1 Introduction
The current scholarship on behavioral responses to mass shootings — an example of a crisis event that legislators might respond to — has particularly focused on the effects of these events on public opinion (Gunn et al., 2018; Barney and Schaffner, 2019; Newman and Hartman, 2019; Rogowski and Tucker, 2019), the electoral behavior of affected constituencies (Hassell, Holbein and Baldwin, 2020), and the firearm-related policy environments of states (Luca, Malhotra and Poliquin, 2020), but there are mixed results regarding what effects these events may have on a variety of outcomes. In some instances, scholars find that mass shootings cause surges in engagement with gun policy (Reny et al., 2023) while others find that responses are limited by temporal, spatial, and political factors (Sharkey and Shen, 2021). In a similar vein, it appears that some mass shootings — particularly those with high levels of media coverage — coincide with increases in firearms purchasing while others do not (Iwama and McDevitt, 2021; Liu and Wiebe, 2019). Additional evidence suggests that legislators may lack the ability or incentive to adjust their behavior even when facing possible ideological, environmental, or electoral pressures (Butler and Hassell, 2018; Bromley-Trujillo and Poe, 2020). Even if we believe that mass shootings might lead to shifts in opinions on firearm-related policies — which is not obvious, as per Baxter-King (2024) — there may not be enough of a change in the electoral behavior of affected constituencies to cause a legislator to alter their voting behavior on relevant policies, even as some scholars find that members of Congress who are facing reelection pressures (and represent sizable “pro-gun constituencies”) may be more likely to change how they vote on gun-related legislation (Bouton et al., 2014). Furthermore, it is not clear that changes in the voting behaviors or policy preferences of a legislator’s constituency (should they occur) would prompt changes in a representative’s actions (Rogers, 2017).
@@
Our paper investigates the effects of exposure to mass shootings on the voting behavior of state legislators who represent the communities most directly impacted by these events. We focus our analysis on estimating the aggregate effect of mass shootings on the behavior of lawmakers rather than shooting-specific effects. Focusing on the average effect boosts our statistical power and precision by pooling rare and noisy events, while also yielding an informative estimate of how legislators typically respond. This approach avoids the pitfalls of multiple comparisons and overfitting to temporally idiosyncratic incidents, and relies on methods that account for heterogeneity across events to identify a single average treatment effect. We combine roll-call voting records of California state legislators with a novel dataset from the Giffords Law Center (GLC) that ranks every firearm-related bill introduced in the California State Assembly and Senate from 2011 to 2022 on a permissive to restrictive scale. [^4] We validated the GLC’s bill rankings, which is discussed in greater detail in A.3. Using pairwise comparison and the Bradley-Terry model (Bradley and Terry, 1952), we construct a yearly score of each individual lawmaker’s directionality of support for gun control legislation from the 2011-2012 to 2021-2022 sessions. Our scoring regime covers 264 individual California legislators over twelve years.
@@
Using this measure of an individual legislator’s latent position on firearms policy over time, we identify which legislators represented districts that experienced a mass shooting between 2011 and 2022. Through a difference-in-differences design accounting for heterogeneity in treatment timing, we estimate whether the occurrence of a mass shooting within a legislator’s district causes a measurable change in their position on gun policy. We find that when a mass shooting occurs within a legislator’s district, California lawmakers reduce their support for gun control policies by approximately 1.15% compared to all other legislators (ATT = -0.023, SE = 0.066, outcome measured on a -1 to 1 scale), though this estimate is a statistically insignificant null. Our results find no evidence of responsiveness in a legislators positionality on firearms-related policy and these statistically insignificant estimates persist across multiple estimation strategies and varied treatment assignments (such as estimations accounting for the distance of a legislator’s district from a mass shooting). This finding conflicts with the self-reported assessments of responsiveness to mass shootings that these same legislators described in a series of qualitative interviews we conducted. In sum, our mixed-methods approach in California finds that legislators do not change their voting behavior following a mass shooting — though this finding only speaks to a single metric of legislator behavior in a single state.
@@
As a robustness check on our main findings for California, we analyze the effects of mass shootings on the firearm policy positions of state legislators in five additional states: Colorado, Florida, Georgia, Missouri, and Texas. [^5] The GLC ranks every state in terms of how restrictive the gun policy environment is. In this paper, we focus on states that had at least one mass shooting occur during our period of study and select one state from each of the quintiles of the GLC ranking: California (1), Colorado (10), Florida (24), Georgia (33), and Missouri (48). Texas (31) is also included as it experienced a large number of shootings (6) during our period of study, and has a more permissive policy environment than California. This offers an opportunity to assess the effects of mass shootings in varying firearms policy environments. These states all experienced at least one mass shooting during the time period we study in this paper. Ultimately, we find that across a twelve year period, 28 mass shootings — occurring in 6 states and requiring the scoring of nearly 1,900 individual state legislators over time — caused no discernible change in the firearm policy positions of state legislators on regulatory proposals on firearms, on average, when such an event occurs within their districts. Exceptions do occur — when we disaggregate our results and estimate shooting-specific effects, we find evidence that the aggregation method we utilize to account for varied treatment timing masks heterogeneity in the treatment effects of individual shootings. However, our null results hold when we also inspect other forms of behavior that legislators may engage in: we find that legislators do not cosponsor or author more or fewer firearm-related bills following a mass shooting, on average.
@@ 2 Theory and Qualitative Findings
To better inform our expectations regarding the effects of mass shootings on the voting behaviors of state legislators, we conducted a series of qualitative interviews with five current and former members of the California state legislature who represented districts in which a mass shooting had occurred between July 2010 and 2022.666Additional details on the procedures and results of these qualitative interviews can be found in section A.1 of the Appendix. In this section, we introduce the literature that informs our theoretical expectations, and discuss the results of these qualitative interviews that support and contradict our priors.
@@ 2.1 Do Mass Shootings Incentivize Agenda Setting or Act as Windows of Policy Opportunity?
Political agenda setting — which can be conceptualized as political elites’ ability to influence public opinion and public policy, for instance — has long been considered the “prime instrument of power” wielded by lawmakers to alter the nature of political conflict (Schattschneider, 1960). At the sub-national level, political agenda setting is largely driven by questions of issue prioritization, as state or local officials determine which areas to focus their rather limited policy making resources on addressing (Franklin and Ebdon, 2004; Erie, Kogan and MacKenzie, 2011). Mass shootings, as highly salient events that capture the attention of the mass public and media, may then serve as an opportunity for state legislators to set legislative priorities and build on the public’s attention towards specific issues. As Legislator A (D, Southern California) said in their interview with us, the shooting in their district offered them an opportunity to “use the moment, in all its horrifying nature, to also give [the victims] lives and their deaths meaning.”
@@ 2.2 The Limited Effects of Mass Shootings in an Era of Partisan Polarization
The results from our qualitative interviews suggest that partisan polarization may play a substantial role in informing the responsiveness of legislators to mass shootings that occur within their respective districts. Legislator E, a Democrat, spoke about how partisan polarization certainly dictated their reaction to the shooting in their district, saying “[I was] trying to walk that very fine line because of political differences in a community that’s more conservative. There were definitely some families [of victims] that were a little more Second Amendment oriented than others.” Legislator C, a Republican, spoke about the role of partisan polarization more bluntly, explaining what their party told them after the shooting: “I was told directly, if I speak about [gun control], it’s going to cost \$1,000,000 per Tweet. For us to defend you in the next election, it’s going to cost 1,000,000 bucks a Tweet.” The direct threat of sanctions from party leadership that would affect a legislator’s campaign warchest suggests that any sort of defection — even one as mundane as a Tweet — from policy positions preferred by the party would result in severe consequences. [^8] We cannot confirm the veracity of Legislator C’s claims about sanctions threatened by the California GOP. We can, however, confirm that Legislator C did not post a single Tweet related to a firearms policy proposal until 181 days after the shooting. This Tweet was expressing opposition to a permissive policy.
@@ 3.1 Mass Shootings in Scope of Study
Note: Red districts are those affected by a mass shooting, shown as the small, light red points at the location of the shooting. The map uses district boundaries adopted after the redistricting cycle following the 2010 Census, and only plots shootings that affected districts under these new maps (accounting for variation in timing of adoption of these maps across states). For a full list of affected districts and shootings (including those happening prior to redistricting), see A.2. Nebraska has a unicameral state legislature, so we do not report lower chamber districts and only show upper chamber districts on the map.
Additional details on these shootings and the selection decisions underlying the inclusion of these states in our analyses can be found in Table A.2.1. For California, we consider the effects of eleven mass shootings.
@@ 3.2.1 GLC Firearm Bill Rankings and Bill Effects
From 2011 to 2023, the GLC hand-coded every bill introduced or enacted by every state legislature that was directly related to firearms policy — with the universe of legislation ranging from provisions allowing the permitless carry of concealed firearms, to large-scale budgetary items mentioning firearms in a single sub-item, to bans on assault-style weapons. This yields a total of 15,299 legislative attempts at the state level. For each of them, the GLC ranked it on a “weaken/neutral/strengthen” scale with regards to the bill’s likely effect on a state’s respective gun control policy environment (with “strengthen” bills making a state policy more restrictive and therefore more favorable to gun violence prevention advocates, for instance). [^10] Under the terms of a non-disclosure agreement dictating the amount and type of data that we are able to share in replication materials and publications, we arranged for conditional access to this as-of-yet untapped source of internal lobbying data.
@@
In section A.3 of the Appendix, we discuss our validation procedure used for the GLC Bill Rankings, and we ultimately accept that “weaken” laws (according to the GLC) can generally be considered permissive (-1), “neutral” as neutral (0), and “strengthen” laws as restrictive (+1). We refer to this -1/0/+1 score as the “Bill Effect” throughout this paper — a novel measure of the relative restrictiveness or permissiveness of a firearm-related legislative measure introduced in a state. Table 1 shows three examples of bills from California during the period that we study, with the original GLC ranking and the Bill Effect.
@@ 3.2.2 Measuring Legislators’ Latent Position on Firearm-related Issues
Assuming that each pairwise comparison is independent, the latent trait $\lambda$ can be estimated by maximum likelihood. As we wish to compare legislators across time, we utilize at least two “bridge” legislators for every state and year to anchor the latent scores ($\lambda_{i}$) to establish intertemporal comparability. We assume that the underlying positionality of these legislators on gun control changes only modestly from one year to the next, and we utilize never-treated chamber leaders (i.e., Speaker, Speaker Pro-Tempore, Senate President Pro-Tempore, etc.) to identify these relatively stable bridging agents. [^12] This assumption does not imply cross-state comparability. See Bailey (2007); Lewis and Tausanovitch (2015); Lewis and Sonnet (2020) for discussions on the limitations on the comparability of ideal point estimates. Finally, we rescale the scores to range from -1 to 1, from least to most supportive of restrictive firearms policies. For California legislators, the Gun Control Scores are calculated using 1,927 votes on firearm-related issues (on average, 160.6 votes each year).
Unlike the conventional methods of using a spatial choice model or Bayesian Item Response Theory (IRT) model on a subset of roll-call voting records, our scoring regime uses the direct Bill Effects to orient scores in the direction of supporting gun control policies. As robustness checks, we also estimate latent traits using the IRT model (Clinton, Jackman and Rivers, 2004) in A.6.1 and discuss alternative measures such as survey responses in A.4.1. Figure 2 shows the distribution of the Gun Control Scores over time for California state legislators. We pick up a small, upward trend for all California state legislators over the years, indicating that California’s gun policy environment has become more restrictive just as the majority of legislators have become more favorable to restrictive policies, on average. However, there is still a clear separation between Democratic and Republican legislators, suggesting that partisan legislators are polarized on this policy domain.
@@ 3.3 Treatment Assignment and Causal Identification
As seen in Figure 1 and Table A.2.1, the shootings we consider as “treatments” in this paper happened in different legislative districts across a wide time period, and in different states. Given this heterogeneity in treatment timing and location, traditional difference-in-differences (DiD) estimators (such as two-way fixed effects) are not appropriate for this setting as they fail to account for fluctuations in the effects of treatment due to adaptation, changing environmental conditions, or the immediacy of responses among treated units (Goodman-Bacon, 2021). For these reasons, we follow a stacked DiD estimation strategy similar to one used by Cengiz et al. (2019) to account for varied treatment timing and treatment effects. In Section A.5, we validate these results by employing the staggered DiD estimation strategy proposed by Callaway and Sant’Anna (2021).
@@
The two central assumptions underlying these approaches are that of heterogeneous treatment adoption and parallel trends with never-treated units. For this first assumption, we posit that once units (legislators in state A) receive a treatment (a shooting in state A occurring within or near their district, depending on the specification), they remain treated throughout the entire period of observation and treatment is irreversible. For parallel trends with never-treated units, we assume that groups of units during periods where treatment is not yet applied serve as “clean” control units — in other words, legislators who have not had a mass shooting occur within or near their district (and never do within the time frame we study) serve as controls — and these units would have followed similar trends in voting behavior in all post-treatment periods. This assumption relies on two conditions that our setting satisfies: (1) there must be a sufficiently powered group of units that never received treatment (we have over 200 never-treated legislators in California, for example), and (2) these never treated units must be similar to those that receive treatment so that we can compare their outcomes. Additionally, we assume that there is limited anticipation of the treatment among all groups of units that are eventually treated. We report our validation and sensitivity analysis for these assumptions in Section A.5.3.
3.3.1 Formal Model
For our stacked DiD estimation strategy ala Cengiz et al. (2019), we create shooting-specific 2 $\times$ 2 datasets with clean controls for our California Analysis — in other words, datasets (g) including the outcome variable for the treated group and all other observations that are never treated within the observation window, removing units that were previously treated or will be treated in the future. These datasets are then stacked together and, with a shooting-specific identifying variable, a two-way fixed effects (TWFE) DiD regression is estimated on the stacked data, via the following regression
[…]
where $y_{it}$ denotes the Gun Control Score for legislator $i$ in year $t$, ultimately yielding the average treatment effect on the treated units (ATT). Under this specification, we utilize relative-time indicators (i.e., “Years Since Shooting”) for treatment ($D_{it}^{l}$) instead of a binary treatment indicator (i.e., $D_{it}$) from a canonical TWFE DiD regression. Here, $D_{it}^{l}$ is an indicator for legislator i in year t and group $E_{i}$ being k periods from the start of treatment (i.e., $D_{it}^{l}=\mathbb{I}[t-E_{i}=k]$), such that the first summation captures time periods leading to treatment, and the second summation captures those following treatment.
[…]